OpenStudy (anonymous):

WILL AWARD MEDALS Eliminate the parameter. x = t2 + 2, y = t2 - 4 Options: a. y = x - 6, x ≥ 1 b. y = x + 6, x ≥ 1 c. y = x2- 6, x ≥ 1 d. y = x2+ 6, x ≥ 1

OpenStudy (anonymous):

I don't think it's b

OpenStudy (anonymous):

Yeah, it's not b

ganeshie8 (ganeshie8):

subtract them both

ganeshie8 (ganeshie8):

x = t2 + 2, y = t2 - 4 ---------------- x-y = ?

OpenStudy (anonymous):

-6?

ganeshie8 (ganeshie8):

x = t2 + 2, y = t2 - 4 ---------------- x-y = 0 + 6 x-y = 6 y = x - 6

OpenStudy (anonymous):

So it's going to be y = x - 6, x ≥ 1

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

Thank you! @ganeshie8